OpenStudy (anonymous):

ted drove from boston massaachusettts to buffalo, new york, a distance of 616 miles. he stopped for food and gad, and his breaks took 2 hours. if the trip took 16 hours altogether, what was his average speed? a. 55 miles per hour b. 5 miles per hour c. 44 miles per hour d. 66 miles per hour

OpenStudy (anonymous):

16 hrs subtract the 3 2 hours for food,gas, and breaks =14 hours 616 miles divided by 14 hours = 44 miles per hour

OpenStudy (richyw):

average speed:\[\frac{616}{16}\]average speed while driving:\[\frac{616}{16-2}\]

OpenStudy (richyw):

the first one is not a multiple choice answer, so I would go with the later. Although to me the first one is "more correct"

OpenStudy (whpalmer4):

I agree that his average speed for the trip is 616/16=38.5, not 616/14=44 as the problem seems to want.